OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (1, -5) (5, -1)

OpenStudy (anonymous):

First find m (the slope) by subtracting the 1st point from the 2nd and then use point one to fill in for y1 and x1 when putting it in the form y=mx+b

OpenStudy (triciaal):

slope,m = change in y /change in x = (y2 -y1)/(x2 - x1) given two points on the line (x1, y1) and (x2, y2) using either of the points and using y = mx + b find b write the equation of the line y = mx + b for any point (x, y)

OpenStudy (triciaal):

do you need me to check your work?

OpenStudy (anonymous):

y=-1x+4?

OpenStudy (triciaal):

m = -5--1/1-5= 1

OpenStudy (triciaal):

no

OpenStudy (triciaal):

-5 = 1(1) + b -5 -1 = b = -6 y = x -6 check with point 2 -1 = 5 -6 = -1

OpenStudy (triciaal):

got it?

OpenStudy (triciaal):

you had calculated the slope wrong. try and make it a practice to check your work. with your equation y = -x + 4 using (1, -5) y = -5 -x + 4 = -1 + 4= 3 -5 not = 3 so you would see you have an error